A clothing company hires workers in the competitive market to cut denim to make jeans. The fabric costs $5. The weekly output of the company varies with labour usage and is shown below: Number of worke Jeans (pairs/week) 20 35 48 60 a) What is the marginal productivity of each worker? (1p) b) If the jeans sell for 35 a pair, what is the value of the marginal product of each c) If the competitive market wage is 400 per week, how many workers should the d) worker? (2p) company hire? Explain with a maximum of 1 sentence to why that is the case (1p). How many pairs of jeans will the company produce each week? Explain with a maximum of 1 sentence to why that is the case (Ip)
Answer #1
No of workers Jeans(pairs/week) MP of workers Price of product Value of MP(€)
0 0 0 €35/pair 0
1 20 20-0=20 35 35*20=700
2 35 35-20=15 35 35*15=525
3 48 48-35=13 35 35*13=455
4 60 60-48=12 35 35*12=420

C) If the competitive market wage rate is €400/week then firm should use 4 workers. Because firm will use the labour as long as value of marginal product is greater than wage paid to the workers. In 4 units of labour use value of marginal product of labour is €420 and it is more than€400.

D) The company would produce total 60 pairs of jeans per week. Because the company will use 4 workers.